OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. Y=3/4x-9; (-8,-18)

OpenStudy (anonymous):

do you know point-slope form?

OpenStudy (anonymous):

if it is paralell to the other line then slope is same as that which is from given eq is=3/4 so now you have one point (-8,-18) and slope=m=3/4 now use point slope form that is.. y – y1 = m(x – x1) use this and solve...